28 February 2012

Cloudy with a chance of Maple Leaves

The temperature in Eureka was a cold -47 °C this morning, and rose to -41 °C in the afternoon. Thanks to the temperature inversion, PEARL enjoyed somewhat warmer -30 °C. A thick layer of clouds blocked the Sun for the entire day, preventing Bruker and PARIS measurements. Joseph helped Xiaoyi by finishing and analyzing resolution tests for the UT-GBS. Debora ran her Linefit program using HBr cell measurements taken earlier in the campaign to check the calibration of the instrument. She compared them to previous years' results to check that the spectra taken by PARIS are in a good shape. Debora and Volodya also attempted to revive a damaged network card, used for the internet connection used for data backups, and created a temporary fix.

Early this afternoon, the team took a break to raise a new Canadian flag on the PEARL flagpole. The old flag had become worn from the Arctic climate. It was the first time some of us had raised our national symbol, and was a powerfully uplifting experience.

Xiaoyi has continued his work with the PEARL-GBS in the UV/Vis lab. Last night, it again stopped taking dark current measurements prematurely. Fortunately, Xiaoyi appears to have now isolated the issue to the instrument computer's communication with the sun tracker. This problem should be solved with relative ease by adjusting the labview code. Having consulted with an ACE Campaign and GBS veteran, Cristen, Xiaoyi again left the PEARL-GBS taking dark current measurements overnight with the sun tracker disabled. He feels confident tomorrow will finally yield success!

Meanwhile, Pierre worked on the roof to prepare for the installation of the new sun tracker. If the weather is good, he and Xiaoyi will install the sun tracker tomorrow.

Also in the UV/Vis lab today, Volodya tested Brewer 021. The SPS and MAESTRO continue normal operation collecting zenith sky data, which will be used to retrieve over 20 different atmospheric species.

During the day, Emily ran several calibration runs for the CRL lidar system at 0PAL, two with white light and two with bright skylight. She also installed the sheet polarizer in the depolarization channel. In the evening, Chris and Emily worked together to coordinate optimizing the photo-multiplier tube focus and the rotation angle of the polarizer to maximize the depolarization channel signal. We now receive sufficient photons for this channel to be useful for measurements. With this modification complete, Chris ran the lidar for 10 hours overnight, taking data in all receiver channels.

The first ACE ozonesonde flight successfully occurred today. The balloon was launched at 18:05 and reached an altitude of 33,500 m and 5.1 millibar.
The Polar Vortex continues to be centered over the Arctic Ocean. The edge stretches from northern Russia to the northernmost edge of Ellesmere Island.

Also in the news today, the end of full-time year-round operations at PEARL as of April 30th was formally announced. The closure notice can be found here. The French version can be found here.

Dan
Back to Daily Reports
Back to Home